OpenStudy (anonymous):

calculate the new freezing point of the substance if 292 g of calcium nitrate is added to 50.0 L of the substance. The density of the pure substance is 1.16 g/mL. It is a polar substance, and its molal freezing point constant, Kf is .450 def C/m pls help, not sure how to start this problemm

OpenStudy (anonymous):

do you know the formula to calculate the freezing point depression?

OpenStudy (anonymous):

are you still around?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the formula is delta Tf = Kf * dissociation factor * molality

OpenStudy (anonymous):

yup..! good. but here, the solute is calcium nitrate. which dissociates into two Ca^2+ and NO3 ^2- so, you will multiply Kf and molality by 2 which is the van't hoff factor you need to calculate the molality now. to find the delta Tf. \[molality = \frac{ no. of moles of solute }{ weight of solvent \in kilograms. }\] can you calculate the molality now? to get the weight of solvent, you are provided with the volume of the solvent and it's density too. i think you can now calculate the weight of solvent and put all the data in the formula and get delta Tf.

OpenStudy (anonymous):

wait the dissociation factor should be 3 right? calcium nitrate is Ca(NO3)2

OpenStudy (anonymous):

ohh yes yes.. my mistake. its 3 not 2 you r right!

OpenStudy (anonymous):

ok so moles of solute is 1.78 moles

OpenStudy (anonymous):

but this is the part idk how to do. how do i find the weight of the solvent in kilograms?

OpenStudy (anonymous):

\[Density = \frac{ Mass }{ Volume }\] density = 1.16 g/mL Volume of solvent = 50L = 50,000mL now can you find the mass.? the answer will come out in grams.

OpenStudy (anonymous):

sounds sense??

OpenStudy (anonymous):

ok so mass = density * volume 1.16 g/ml * 50,000 mL = 58000 g

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

then i take 1/78 moles/ 58 kg = .0307 m

OpenStudy (anonymous):

1.78*

OpenStudy (anonymous):

then i plugged it into the equation and got .0414 as delta Tf

OpenStudy (anonymous):

should i subtract tht from .450 degrees? because it is asking for the new freezing point

OpenStudy (anonymous):

is 450 your original freezing point ?? i.e. freezing point of solvent??

OpenStudy (anonymous):

yes i think soo. it says the freezing point constant is .450

OpenStudy (anonymous):

if so, then yes. you must subtract the delta Tf from 450. make sure that both are in same units. i guess the delta Tf is in celcius. for the unit of Kf is in celcius. make sure that it's 450C and if this is the case, just subtract and get the new freezing point

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thank you so muchh!!

OpenStudy (anonymous):

welcome!